Problem 6 - Based on the data provided in the table below is this new forecasting technique biased? If so how is it biased? Calculate tracking signal values. Is there any evidence this New Forecast should not be used?

Actual New
Period Demand Forecast
1 650 610
2 883 932
3 950 917
4 1025 988
5 680 853
6 895 788
7 988 942
8 1178 1083
9 725 952
10 915 820
11 1068 992
12 1256 1162
Answer #1
Period Actual Demand New Forecast Forecast Error = Actual - Forecast Running Sum of Forecast error (Keep adding forecast errors) Square of error Absolute deviation
1 650 610 40 40 1600 40
2 883 932 -49 -9 2401 49
3 950 917 33 24 1089 33
4 1025 988 37 61 1369 37
5 680 853 -173 -112 29929 173
6 895 788 107 -5 11449 107
7 988 942 46 41 2116 46
8 1178 1083 95 136 9025 95
9 725 952 -227 -91 51529 227
10 915 820 95 4 9025 95
11 1068 992 76 80 5776 76
12 1256 1162 94 174 8836 94
# of Observations 12 Total 134144 1072
MSE = 134144/12 11179
MAD = 1072/12 89.33
Tracking Signal = RSFE / MAD = 174/89.33 1.95 Underestimating the demand

MSE = SUm of Mean square errors / # of Observations

MAD = Sum of absolute deviation / # of Observations

Tracking Signal =RSFE / MAD = 1.95. A positive tracking signal indicates that the forecast has a bias and is underestimating the demand consistently i.e. demand is greater than forecast. Hence, it is not a good forecasting method.

Further, the Mean square error is quite high and the forecasting technique is not good.